OpenStudy (anonymous):

Jeremy is going to help his neighbor build a patio similar in shape to his patio. Jeremy's neighbor's patio will have twice the area of Jeremy's patio. How can Jeremy use the dimensions of his patio to determine the dimensions of his neighbor’s patio? A. Multiply each dimension by 2004-02-04-06-00_files/i0220001.jpg. B. Double each dimension. C. Square each dimension. D. Add 5 to each dimension.

OpenStudy (anonymous):

Option A is Multiply by the square root of 2.

OpenStudy (misssunshinexxoxo):

The figure are similar, corresponding sides are proprtional L1/L'1=L2/L'2= ... =k , the scale factor for lengths. However, if scale factor for lengths is k, the ratio of area is k^2 find Scale factor Use the given area to find the proportionality factor then use it the ratio of length to find the missing distance between the parallel sides. Area of new /Area of old=360/250=36/25=k^2 k=SQRT(36/25)=6/5 Ratio of lengths =k=6/5 Distance between parallel sides of new patio/Distance between parallel sides of old patio=6/5 Distance between parallel sides of new patio= (6/5)*12.5=15

OpenStudy (anonymous):

Thank you but that does not answer the question..

OpenStudy (anonymous):

How can Jeremy use the dimensions of his patio to determine the dimensions of his neighbor’s patio? A. Multiply each dimension by 2004-02-04-06-00_files/i0220001.jpg. B. Double each dimension. C. Square each dimension. D. Add 5 to each dimension.